# M3-JEPA: Multimodal Alignment via Multi-gate MoE based on the Joint-Predictive Embedding Architecture

### v5

@@ Abstract
Current multimodal learning strategies primarily optimize in the original token space. Such a framework is easy to incorporate with the backbone of pretrained language model, but might result in modality collapse. To alleviate such issues, we leverage the joint embedding predictive architecture (JEPA) on the multimodal tasks, which converts the input embedding into the output embedding space by a predictor and then conducts the cross-modal alignment on the latent space. We implement this predictor by a Multi-Gate Mixture of Experts (MMoE) and name the framework as M3-JEPA, accordingly. The gating function disentangles the modality-specific and shared information and derives information-theoretic optimality. The framework is implemented with both contrastive and regularization loss, and solved by alternative gradient descent (AGD) between different multimodal tasks. By thoroughly designed experiments, we show that M3-JEPA can obtain state-of-the-art performance on different modalities and tasks, generalize to unseen datasets and domains, and is computationally efficient in both training and inference. Our observation suggests that M3-JEPA might become a new basis to self-supervised learning in the open world.
@@ 1 Introduction
Although these studies achieve remarkable success in multimodal learning, however, they may be often subject to the issue of modality collapse during the cross-modal alignment, resulting from the conflicting gradients (Javaloy et al., 2022), missing modality or labels (Wu et al., 2024b), and mismatched data distribution or fusion (Ma et al., 2022). Especially, LLMs take advantage of self-supervised learning (SSL), which predicts the hidden parts of the data grounded by the visible parts, in the discrete token space. Nevertheless, this paradigm relatively struggles in continuous domains (e.g., image or video), which causes the cross-modal alignment difficult to converge, and might fail to capture key information (Dawid & LeCun, 2024), especially in the existence of information uncertainty, redundancy, or ambiguity (e.g., one picture can be described by two content different but semantically similar sentence).
To mitigate these issues, one possible solution is to switch from generative and probabilistic modeling to the energy-based model (EBM) paradigm, which minimizes a non-negative energy function of both the input and output embeddings. To better handle the uncertainty of information, the joint embedding predictive architecture (JEPA) (Dawid & LeCun, 2024) is proposed, which employs a latent predictor that projects the necessary information from the input into the output embedding space, while filtering out irrelevant or misleading input features. JEPA then aligns the input and output signals in the latent level instead of the token or pixel level, with a similar idea proposed by the ”platonic representation” (Huh et al., 2024). JEPA has been applied in vision, namely I-JEPA (Assran et al., 2023), in which an arbitrary part of an image is masked and predicted by the other image blocks; as well as the motion and content features of videos (Bardes et al., 2023). Nevertheless, there has been no generalized framework that applies the idea of JEPA to generalized multi-modality modeling, with any-to-any modality combinations either for the input or the output.
@@ 2 Preliminary
2.1 Joint-Predictive Embedding Architecture
@@ 4.1 The Information-Theoretic Analysis
The above statements ensure that minimizing the averaged task-specific losses ($\mathcal{F}(x,y),\mathcal{F}(y,x)$) in Equation 11 is equivalent to the information-theoretic objective in Equation 15. Also, the loss weight $\alpha$ in both equations is consistent, balancing the compression ($\mathcal{L}_{cl}$ and $\mathcal{I}$) and predictiveness ($\mathcal{L}_{reg}$ and $\mathcal{H}$).
@@ 5.1 Setting
All tasks have the batch size of 128, solved by the Adam optimizer with the lr schedule of cosine, warmup of 0.1 and weight decay of 0.005. For retrieval tasks, we evaluate recall-based metrics including R@1, R@5 and R@10. For classification tasks, we provide metrics such as Accuracy, Precision, Recall and F1 score. Further details of implementation, datasets and benchmarks are summarized in Appendix LABEL:sec:_implementation.
@@ 6.2 Multimodal Models with MoE Structures
There are also pioneer works that integrate multimodal models with MoE-like structures. For instance, LIMoE (Mustafa et al., 2022) feeds both images and text signals into a single MoE encoder, trained by a contrastive loss. VL-MoE (Shen et al., 2023) uses modality-specific experts for image-text modeling. MoE-LLaVA (Lin et al., 2024) proposes MoE-Tuning, a training strategy for large vision-language models, and construct a sparse model with a constant computational cost. In contrast, M3-JEPA adopts the multi-gate MoE architecture, which simultaneously feeds parallel output terms to both contrastive and regularization losses.
@@ A.3 Image Classification
We train and test M3-JEPA on the popular benchmark, ImageNet-1K (Deng et al., 2009), which contain 1,281,167 training images, 50,000 validation images, and 100,000 test images across totally 1,000 classes. For image classification, the initial learning rate is 1.0e-3 and the final learning rate is 5.5e-6.

### v6

# M3-JEPA: Multimodal Alignment via Multi-gate MoE based on the Joint-Embedding Predictive Architecture

@@ Abstract
Current multimodal learning strategies primarily optimize in the original token space. Such a framework is easy to incorporate with the backbone of pretrained language model, but might result in modality collapse. To alleviate such issues, we leverage the Joint-Embedding Predictive Architecture (JEPA) on the multimodal tasks, which converts the input embedding into the output embedding space by a predictor and then conducts the cross-modal alignment on the latent space. We implement this predictor by a Multi-Gate Mixture of Experts (MMoE) and name the framework as M3-JEPA, accordingly. The gating function disentangles the modality-specific and shared information and derives information-theoretic optimality. The framework is implemented with both contrastive and regularization loss, and solved by alternative gradient descent (AGD) between different multimodal tasks. By thoroughly designed experiments, we show that M3-JEPA can obtain state-of-the-art performance on different modalities and tasks, generalize to unseen datasets and domains, and is computationally efficient in both training and inference. Our observation suggests that M3-JEPA might become a new basis to self-supervised learning in the open world.
@@ 1 Introduction
Although these studies achieve remarkable success in multimodal learning, however, they may often be subject to the issue of modality collapse during the cross-modal alignment, resulting from the conflicting gradients (Javaloy et al., 2022), missing modality or labels (Wu et al., 2024b), and mismatched data distribution or fusion (Ma et al., 2022). Especially, LLMs take advantage of self-supervised learning (SSL), which predicts the hidden parts of the data grounded by the visible parts, in the discrete token space. Nevertheless, this paradigm relatively struggles in continuous domains (e.g., image or video), which causes the cross-modal alignment difficult to converge, and might fail to capture key information (Dawid & LeCun, 2024), especially in the existence of information uncertainty, redundancy, or ambiguity (e.g., one picture can be described by two content different but semantically similar sentence).
To mitigate these issues, one possible solution is to switch from generative and probabilistic modeling to the energy-based model (EBM) paradigm, which minimizes a non-negative energy function of both the input and output embeddings. To better handle the uncertainty of information, the Joint-Embedding Predictive Architecture (JEPA) (Dawid & LeCun, 2024) is proposed, which employs a latent predictor that projects the necessary information from the input into the output embedding space, while filtering out irrelevant or misleading input features. JEPA then aligns the input and output signals in the latent level instead of the token or pixel level, with a similar idea proposed by the ”platonic representation” (Huh et al., 2024). JEPA has been applied in vision, namely I-JEPA (Assran et al., 2023), in which an arbitrary part of an image is masked and predicted by the other image blocks; as well as the motion and content features of videos (Bardes et al., 2023). Nevertheless, there has been no generalized framework that applies the idea of JEPA to generalized multi-modality modeling, with any-to-any modality combinations either for the input or the output.
@@ 2 Preliminary
2.1 Joint-Embedding Predictive Architecture
@@ 4.1 The Information-Theoretic Analysis
The above statements ensure that the averaged task-specific losses ($<\mathcal{F}(x,y)>$, specified by Equation 8-11) is equivalent to the information-theoretic objective in Equation 15. Also, the loss weight $\alpha$ in both equations is consistent, balancing the compression ($\mathcal{L}_{cl}$ and $\mathcal{I}$) and predictiveness ($\mathcal{L}_{reg}$ and $\mathcal{H}$).
@@ 5.1 Setting
All tasks have the batch size of 128, solved by the Adam optimizer with the lr schedule of cosine, warmup of 0.1 and weight decay of 0.005. For retrieval tasks, we evaluate recall-based metrics including R@1, R@5 and R@10. For classification tasks, we provide metrics such as Accuracy, Precision, Recall and F1 score. Further details of implementation, datasets and benchmarks are summarized in Appendix A.
@@ 6.2 Multimodal Models with MoE Structures
There are also pioneer works that integrate multimodal models with MoE-like structures. For instance, LIMoE (Mustafa et al., 2022) feeds both images and text signals into a single MoE encoder, trained by a contrastive loss. VL-MoE (Shen et al., 2023) uses modality-specific experts for image-text modeling. MoE-LLaVA (Lin et al., 2024) proposes MoE-Tuning, a training strategy for large vision-language models, and constructs a sparse model with a constant computational cost. In contrast, M3-JEPA adopts the multi-gate MoE architecture, which simultaneously feeds parallel output terms to both contrastive and regularization losses.
@@ A.3 Image Classification
We train and test M3-JEPA on the popular benchmark, ImageNet-1K (Deng et al., 2009), which contains 1,281,167 training images, 50,000 validation images, and 100,000 test images across totally 1,000 classes. For image classification, the initial learning rate is 1.0e-3 and the final learning rate is 5.5e-6.